EVA Report – December 13th

Crew 200 EVA Report 12-13-2018

EVA #6

Author: Lindsay Rutter

Purpose of EVA: 1) Map out locations that allow radio communication
between EVA crew and HabCom around Phobos Peak 2) Test how EVAs can
safely and effectively be conducted with three-member crew, including
rover use.

Start time: 14:12

End time: 16:00

Narrative: With media guests visiting the habitat today, our EVA was
pushed back in time, which was authorized by Shannon. We continued our
communication and radio strength and connectivity testing protocol
today. Our EVA began with an engineering check. We then used
Opportunity to drive north on Cow Dung Road. We intended to stop at
the west end of Stream Bed Connector, but we inadvertently stopped at
the east end of Galileo Road 1104, which we walked along for almost
its entirety. Through our GPS system and our communications with
HabCom, we realized we were probably on Galileo Road 1104 and so we
returned back to the west end of it. We then continued driving
Opportunity northward. We intended to explore areas about 1 kilometer
north of Reservoir Dam, but by the time we were 0.2 kilometers north
of Reservoir Dam, we had lost connectivity with HabCom for several
consecutive radio tests. This lack of connectivity continued even when
we parked the rover and stood on higher points. We decided to return
driving south, communicating with HabCom throughout the return,
sometimes having to park the rover and stand on higher points.
Arriving back at the habitat about fifteen minutes early, we explored
the hills on the west side of the habitat, still following our
communication and radio strength and connectivity testing protocol. We
are excited to report that the battery usage on Opportunity was much
more efficient than it has been lately.

Destination: Galileo Road 1104, above 0.2 kilometers north of Reservoir Dam

Coordinates (use UTM NAD27 CONUS): 519100 E – 4252000 N, 518250 E – 4253000 N

Participants: Makiah, Lindsay

Road(s) and routes per MDRS Map : Driving on Cow Dung Road and walking
on Galileo Road 1104

Mode of travel: Driving and walking

Greenhab Report – December 13th

Crew 200 Greenhab Report 13-DEC-2018
Greenhab Officer:
Makiah Eustice

Environmental Control:
40% Shade Cloth on

Average Temperatures:
Low: 24 C
High: 35 C

Hours of Supplemental Light:

Daily Water Usage of Crops:
8 gallons

Water in Blue Tank (lbs):

Times of Watering for Crops:

Changes to crops:

The weird 54 C reading was a trick of the light, literally! I moved it
closer to the center of the hab, away from direct sun. I also tested
with an analog thermometer.


Support/Supplies Needed:
It would be interesting to sees the O2 ans CO2 levels throughout the
day. It would be an important measure to have on Mars! If you can, I’d
love suggestions for sensors.

Sol Summary – December 13th

Sol: 11

Summary Title: The first 3 crew EVA on Mars

Author’s name: Makiah Eustice

Mission Status: Crew conducted an EVA to the North of Stream Bed connector.

Sol Activity Summary: Crew worked on projects in the morning, waiting
to go on EVA after a visit

from guests. We enjoyed home-made bread from Antoine throughout the
day. We made the decision to

end out sim tomorrow, which mean we are wrapping up our projects and
making grand last meals.

Look Ahead Plan: Tomorrow we will wrap up our projects and end
simulation after lunch.

Anomalies in work: None

Weather: Cold, but drier overall.

Crew Physical Status: No crew members reported physical problems today.

EVA: EVA Crew went by Rover to Galileo road, and then just past
Reservoir Dam to investigate

comms capability.

Reports to be file: Operations Report, EVA Report, GreenHab Report,
Pictures Report, Journalist

Report, Science Report

Support Requested: N/A

Journalist Report – December 13th

Sol: 11
Acting Journalist: Lindsay Rutter

Mesas surround us with their red striations. Iron oxide that pumped
through the hearts of long-deceased stars. It is only skin-deep, our
crew geologist says. Dig a few inches and the soil is an anemic grey.

We wake up one morning to the aftermath of heavy snowfall. Our world
radiates a blinding white from all directions and our loud orange
spacesuits only add to the dizzying effect.

Nighttime at last. Our eyes can heal. A jet-black globe encircles us.
We were swallowed by a giant galactic beast. Shy stars I have never
seen before come out from hiding. Light does not litter this isolated
land. We can perceive stars with apparent magnitudes reaching the
limits of human visual perception from this area, our crew astronomer
says. How much deeper into the night sky can other creatures see from
this planet? Their lenses hard at work focusing the light like tiny
organic telescopes.

A spacesuit greets me with a dramatic new makeover as I enter the
habitat one afternoon. A crew member is painting her spacesuit head to
toe with splashes of color. Extravehicular activities are hard work,
she says. The colors represent the energy flowing through different
parts of her body. These senses can be transcribed into color. I ask
her if she has heard of synesthesia.

Entering the GreenHab and I can almost smell the color green. The
plants absorb red and blue and reflect a green perfume. Colors feed
the plants and the plants feed us.

Our crew engineer troubleshoots the power control board, our lifeline
on this harsh terrain. Five lights blink furiously. Red. Green. Red.
Red. Green. He asks me to confirm which color is which. Not everyone
perceives these colors the same.

The Martian flag hangs proudly in the airlock. An unofficial tricolor
design with red, green, and blue stripes. For many, these colors
represents the terraformation of Mars, a topic that remains a matter
of debate. I have experienced many new things during this mission and
have grown from a full spectrum of viewpoints of my crew mates. Mars
means different things to different people, from the red dot to the
blue dot.

Lindsay Rutter

Operations Report December 13th

Crew 200 Operations Report 13-Dec-2018

SOL: 11

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5:30 pm last night to around 9 am today (SOC was still at 68% this morning). On tonight at 4:50pm.

Solar— SOC 77%
Diesel Reading – 100 % (refilled)
Propane Reading – 27 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 340 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 3

Toilet tank emptied: yes

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: yes (by media)
Hours: 43
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: yes (by media)
Hours: 62.1
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: yes (by media and us)
Hours: 42.9
Beginning charge: 100%
Ending charge: 72%
Currently charging: yes

Curiosity rover used: yes (by media)
Hours: 58.8
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were used today by the media, we used Opportunity for our planned EVA, whose batteries did pretty well.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: We repaired a tarp this morning, on the RAM tunnel. No issues to report with the heater.

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this afternoon (we had to delay it as the media were recording this morning, approved by Shannon. Spacesuit 3 still not functional. On the one piece suits, it may be useful to have a chine guard to prevent the chine to hurt the bottom of the helmet especially when driving (mostly for short people).

Summary of Hab operations: repaired a tarp.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report, will be cleaned up.

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site